Joel lifts his spare anvil of mass 75 kg from the floor to a height of 1.3 m onto a bench. 

a. How much gravitational energy does this create for the anvil?

b. If the anvil then falls off the bench what speed will it have when it lands back on the floor? (conservation of energy)


1
Expert's answer
2021-06-07T15:01:27-0400

a)


"U=mgh\\U=(75)(9.8)(1.3)=955.5\ J"

b)


"mgh=0.5mv^2\\v^2=2(9.8)(1.3)\\v=5.05\frac{m}{s}"